OpenStudy (anonymous):

You need a 75% alcohol solution. On hand, you have a 250 mL of a 20% alcohol mixture. How much pure alcohol will you need to add to obtain the desired solution? You will need mL_____ of pure alcohol to obtain mL______ of the desired 75% solution.

OpenStudy (anonymous):

If anyone can explain in steps that would be great, thanks!

OpenStudy (kropot72):

Let the xtra amount of alcohol needed = x (mL) The amount of alcohol in the original solution = 250 * 0.2 = 50 mL The final total amount of 75% alcohol solution = (250 + x) mL The amount of alcohol in the 75% alcohol solution = (50 + x) mL \[0.75= \frac{50+x}{250+x}\] Now you need to solve for x.

OpenStudy (anonymous):

Thank you!

OpenStudy (kropot72):

You're welcome :) Come back if you need more help on this one.

OpenStudy (kropot72):

0.75(250 + x) = 50 + x 187.5 + 0.75x = 50 + x 0.25x = (187.5 - 50) = 135.5 \[x=\frac{135.5}{0.25}\]

OpenStudy (anonymous):

Thank you :)